25. .. Two vectors are given by A = 4.0 x-2.0 and B = 1.0 (c) a vector 5.0 ỷ. Whatis (a)A-B, (b) B-A, and such that + B + = 0?
Answer #1

A+ B = (4hat{x}-2hat{y})+(1hat{x}+5hat{y})
A+ B = (5hat{x}+3hat{y})

Similarly,

B-A = (1hat{x}+5hat{y})-(4hat{x}-2hat{y})
B-A = (-3hat{x}+7hat{y})

Vector C
A+B+C = 0
CA- B
C = -(A+B)
C--(52 + 3の .................From Part a
C = (-5hat{x}-3hat{y})